OpenStudy (anonymous):

Find a unit vector with the same direction as v v= 5i+ sqrt11j

OpenStudy (tkhunny):

Have you considered calculating the magnitude of the vector?

OpenStudy (ikram002p):

unit vector need direction , but its leangth is always 1 so unit vector from v denoted by u is given by \(\bar u =\dfrac{\bar v }{|v|}\) s.t \(|v| \) is the length of v

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (tkhunny):

You didn't answer my question. This will explain your query.

OpenStudy (ikram002p):

ok i'll let @tkhunny to continue with u ^

OpenStudy (anonymous):

I was leaning that to ikram002p

OpenStudy (tkhunny):

Can you calculate the magnitude of the vector? This is the important question.

OpenStudy (anonymous):

@tkhunny i have not considered finding the magnitude but the magnitude would be sqrt(5^2+sqrt(11)^2)= sqrt(36) = 6

OpenStudy (tkhunny):

That's it. Divide your vector by that scalar and you will be done.

OpenStudy (anonymous):

so, 5/6,sqrt11/6 ?

OpenStudy (tkhunny):

This is exactly the formulation provided earlier by ikram002p. Good work. You can check. Calculate the magnitude of your new vector, and make sure you get one (1).